At 300K, 36g of glucose present in one litre of its solution has an osmotic pressure of 4.98 bar. If the osmotic pressure of the solution is 1.52 bar at the same temperature, what would be its concentration ?

Grade:12

1 Answers

Saurabh Koranglekar
3 years ago
Here we have given

π1= 4.98

π2= 1.52

C1= 36/180

C2= ? (we have to find)

Now according to van’t hoff equation

Π = CRT

Putting the values in above equation,we get

4.98 = 36/180RT ------------------------1

1.52 = c2RT ------------------------2

Now dividing equation 2 by 1 ,we get

(c2x 180) / 36 = 1.52 / 4.98

or

c2= 0.0061

Therefore concentration of 2nd solution is 0.0061 M